# Yuxia Fu  Zhizhen Zhang11footnotemark: 1  Yuqi Zhang  Zijian Wang  Zi Huang  Yadan Luo

### v1

@@ 5.1 Experimental Setup
- LIBERO [libero] is a comprehensive benchmark consisting of four distinct task suites: Spatial, Object, Goal, and Long. Each suite contains 10 tasks with 50 demonstrations for every task. The benchmark enables the assessment of the robot’s multi-skill competence, including spatial relationships, object interactions, and task-specific objectives.
- LIBERO-Plus [liberoplus] builds upon the original LIBERO benchmark by introducing seven distinct perturbations, as shown in Figure 4. The benchmark comprises 10,030 tasks, providing diverse settings for evaluating model generalization and robustness under shifts.
- RoboTwin 2.0 [robotwin] serves as a cross-embodiment benchmark for dual-arm manipulation. As shown in Figure 5, we select three embodiments and four tasks for a comprehensive assessment of the cross-embodiment cross-skill ability.
@@
$\mathrm{OpenVLA}$ [OpenVLA] - - 7 $\times$ 4 84.7 88.4 79.2 53.7 76.5
$\mathrm{VLA}$-$\mathrm{Adapter}$ [vla-adapter] - - 0.68 $\times$ 4 99.6 99.6 98.2 96.4 98.5
@@
$\mathrm{OpenVLA}$ TA [TA] Vision Backbones 7 $\times$ 4 56.6 58.0 55.6 6.6 44.2
$\mathrm{OpenVLA}$ TA [TA] All 7 0.0 0.0 0.0 0.0 0.0
$\mathrm{OpenVLA}$ TA [TA] + $\mathbf{S}$ All 7 74.2 82.6 68.8 24.0 62.4
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [TA] All 0.68 0.0 0.0 0.0 0.0 0.0
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [TA] + $\mathbf{S}$ All 0.68 0.0 0.0 0.0 0.0 0.0
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [TA] + $\mathbf{S}$ Except $\mathbf{H}^{L\rightarrow L}$ 0.70 50.2 34.6 0.0 7.4 23.1
$\mathrm{MergeVLA}_{\mathrm{EMR}}$ EMR [emr-merging] 96.0 63.2 62.0 40.6 65.5
$\mathrm{MergeVLA_{TSV}}$ TSV [tsvm] + $\mathbf{S}$ 99.4 97.8 74.4 54.8 81.6
$\mathrm{MergeVLA_{KnOTS}}$ KnOTS[knots] + $\mathbf{S}$ 96.8 98.8 84.8 71.4 88.0
$\mathrm{MergeVLA_{TA}}$ TA [TA] + $\mathbf{S}$ Except $\mathbf{H}^{L\rightarrow L}$ 0.70 98.0 98.8 85.4 76.6 89.7
$\mathrm{MergeVLA_{WUDI}}$ WUDI [wudi] + $\mathbf{S}$ 97.6 98.2 85.6 78.2 89.9
$\mathrm{MergeVLA_{TIES}}$ TIES [ties] + $\mathbf{S}$ 94.8 94.6 91.8 79.4 90.2
Implementation Details. Our vision-language backbone is Qwen2.5-0.5B [qwen25]. By default, we set $l=L$, $k_{r}=8$, mask ratio $\lambda=0.6$, merging scaling factor $\alpha=1$. All finetuning are conducted on a single NVIDIA A6000 Ada GPU (48 GB). Other hyperparameters can be found in Appendix.
@@ 5.2 Results on LIBERO
We first reproduce results of OpenVLA [OpenVLA] and VLA-Adapter [vla-adapter] on four tasks, as highlighted in grey in Table 1. OpenVLA performs notably worse than the others, with an average success rate of 76.5% and only 53.7% on the long-horizon LIBERO-Long task. Our MergeVLA, modified from VLA-Adapter to be merge-friendly, achieves comparable fine-tuning performance, indicating that our structural changes preserve the original capability.
We have tried evaluating single-task finetuned models on unseen tasks (e.g., testing a Spatial expert on the Object suite), all methods achieve 0% success across tasks. This clearly indicates that existing VLA models lack cross-skill generalization. The lower part of Table 1 reports the merging results. Directly applying TA [TA] to OpenVLA fails entirely, while merging only the vision and projector components improves to 44.2%. Adding a task-specific mask further raises the average success rate to 62.4%, though performance remains unbalanced. On VLA-Adapter, tight coupling between the VLM and action expert makes merging difficult—even with masking, unless the final action block is excluded. As shown in the blue-highlighted rows, our MergeVLA consistently outperforms other merging baselines. With TIES [ties] and WUDI [wudi], it achieves balanced results across all four tasks (up to 90.2% average success rate), while remaining lightweight and only 6.5% below fine-tuning performance. These results confirm that MergeVLA enables efficient knowledge reuse and strong multi-task performance with a compact VLA model.
@@
$\mathrm{OpenVLA}$ [OpenVLA] 34.8 0.8 23.0 8.1 28.5 3.5 15.2 16.3
$\mathrm{\pi_{0}}$ [pi0] 81.4 13.8 58.8 85.0 68.9 6.9 79.0 56.3
$\mathrm{VLA}$-$\mathrm{Adapter}$ [vla-adapter] 82.2 53.4 75.0 77.9 74.5 44.6 73.0 67.4
$\mathrm{OpenVLA}$-$\mathrm{OFT}$ [openvla-oft] 93.3 56.4 79.5 88.7 74.2 31.9 75.8 71.4
$\mathrm{MergeVLA}$ 95.5 71.0 78.5 99.0 79.5 51.3 85.1 80.0
@@
$\mathrm{MergeVLA}_{\mathrm{TSV}}$ 78.3 55.0 69.5 89.3 74.7 40.1 70.0 68.1
$\mathrm{MergeVLA}_{\mathrm{TA}}$ 86.3 59.3 72.8 92.7 77.6 39.2 75.9 72.0
$\mathrm{MergeVLA}_{\mathrm{TIES}}$ 94.2 58.2 70.2 93.1 75.3 35.6 78.5 72.2
@@ 5.3 Results on LIBERO-Plus
From Table 2, we first observe that when using single-task finetuned checkpoints evaluated on their own tasks, MergeVLA exhibits stronger robustness under various perturbations than existing VLA models. This improvement stems from its architecture design, which preserves the pretrained VLM’s inherent robustness to visual and language perturbations. Secondly, in the model-merging setting, we find that applying different merging methods with MergeVLA maintains similar robustness even under cross-task evaluation. Notably, when using TA and TIES merging, the merged model even surpasses OpenVLA-OFT in the single-task setting, demonstrating that our MergeVLA can effectively transfer and preserve robustness across tasks.
@@ 6 Conclusion
[…]
@@ 8 Algorithm Details
1:Inputs: Task masks $\{\mathbf{S}_{m}\}_{m=1}^{M}$; Expert heads $\{\mathbf{H}_{m}^{\,l\rightarrow L}\}_{m=1}^{M}$; Pretrained VLM weights $\Theta_{0}$; Merged task vector $\tau_{\mathrm{merge}}$; Value projections of the merged action expert at block (l1)(l{-}1): $\mathbf{V}_{\mathrm{T}}^{\,l-1},\;\mathbf{V}_{\mathrm{A}}^{\,l-1}$; Initial observation $(\mathbf{I}_{0}^{v},\mathbf{I}_{0}^{w},L)$
@@
5:   $\mathbf{h}_{\mathrm{T},m}^{\,l-1},\;\mathbf{h}_{\mathrm{A},m}^{\,l-1}=\Theta_{\mathrm{VLM}}^{(m)}(\mathbf{I}_{0}^{v},\mathbf{I}_{0}^{w},L)$ $\triangleright$ Extract (l1)(l{-}1)-th block hidden states
6:   $\mathbf{V}_{\mathrm{T}}^{\,l-1}=\mathbf{L}_{\mathrm{T}}^{\,l-1}\,\mathbf{\Sigma}_{\mathrm{T}}^{\,l-1}\,(\mathbf{R}_{\mathrm{T}}^{\,l-1})^{\top}$
7:   $\mathbf{V}_{\mathrm{A}}^{\,l-1}=\mathbf{L}_{\mathrm{A}}^{\,l-1}\,\mathbf{\Sigma}_{\mathrm{A}}^{\,l-1}\,(\mathbf{R}_{\mathrm{A}}^{\,l-1})^{\top}$
8:   $\mathbf{r}_{\mathrm{T},m}=\big\|\mathbf{P}_{\mathrm{T}}^{\,l-1}\mathbf{h}^{\,l-1}_{\mathrm{A},m}\big\|_{2}$ $\triangleright$ Choose top-$r_{k}$ singular vectors from $\mathbf{R}$ to get $\mathbf{P}$
9:   $\mathbf{r}_{\mathrm{A},m}=\big\|\mathbf{P}_{\mathrm{A}}^{\,l-1}\mathbf{h}^{\,l-1}_{\mathrm{T},m}\big\|_{2}$
@@ 9 Preliminary Investigation on OpenVLA
In the early stage of this work, we explored the feasibility of directly applying existing model merging methods to OpenVLA [OpenVLA], a popular VLA model. OpenVLA consists of three main components: a vision backbone, a projector, and a language model. The language model itself contains 32 transformer blocks followed by a single-layer MLP head (lm_head). We first attempted to merge all components of OpenVLA using Weighted Average and Task Arithmetic methods. However, the merged checkpoint completely failed on all tasks. This was surprising, as OpenVLA is essentially a VLM, and previous studies [adamms, bring_r_to_v, uq_merge] have shown that VLMs can usually be merged successfully. This prompted us to investigate which part of OpenVLA prevents successful merging.
@@ 9.1 Non-mergeable Components in OpenVLA
To locate the source of failure, we decomposed the model into four submodules: $\mathrm{A.}$ the vision backbone; $\mathrm{B.}$ the projector; $\mathrm{C.}$ the language model body (excluding lm_head); and $\mathrm{D.}$ the lm_head. We then merged each submodule separately across the four official LIBERO task checkpoints using the existing merging method Iso-CTS [iso-cts]. Each merged checkpoint was evaluated on 50 trials per subtask. The results are summarized in Table 7, where the gray-highlighted row denotes the single-task fine-tuning performance. From the table, we observe that merging modules $\mathrm{A}$, $\mathrm{B}$, or $\mathrm{D}$ only slightly decreases success rates, whereas merging the language model body ($\mathrm{C}$) causes complete failure on all tasks. This clearly indicates that the language model is the primary source of merging failure.
@@
Table 7: Success rates on the four LIBERO task suites when merging different components of OpenVLA [OpenVLA] using the Iso-CTS [iso-cts] merging method. Each merged checkpoint combines four task-specific models, while unmerged components retain their original weights. During evaluation, each subtask is tested with 50 trials. Gray-highlighted row indicates the success rates of individually fine-tuned models.
@@ 9.2 Progressive Block-wise Merging of the Language Model
Figure 8: Success rate on the LIBERO-Spatial task when progressively merging the first $k$ language model blocks of OpenVLA [OpenVLA] using the Iso-CTS [iso-cts] merging algorithm. Each configuration merges four task-specific checkpoints and is evaluated over 10 trials per subtask.
@@
Figure 9: Mask active ratios for each LIBERO task suite, computed for both the vision backbone and the language model components following the same definition as in the main paper. Masks are obtained using the Task Arithmetic [TA] merging method with $\lambda=0.6$.

### v2

@@ 5.1 Experimental Setup
- LIBERO [29] is a comprehensive benchmark consisting of four distinct task suites: Spatial, Object, Goal, and Long. Each suite contains 10 tasks with 50 demonstrations for every task. The benchmark enables the assessment of the robot’s multi-skill competence, including spatial relationships, object interactions, and task-specific objectives.
- LIBERO-Plus [15] builds upon the original LIBERO benchmark by introducing seven distinct perturbations, as shown in Figure 4. The benchmark comprises 10,030 tasks, providing diverse settings for evaluating model generalization and robustness under shifts.
- RoboTwin 2.0 [7] serves as a cross-embodiment benchmark for dual-arm manipulation. As shown in Figure 5, we select three embodiments and four tasks for a comprehensive assessment of the cross-embodiment cross-skill ability.
@@
$\mathrm{OpenVLA}$ [28] - - 7 $\times$ 4 84.7 88.4 79.2 53.7 76.5
$\mathrm{VLA}$-$\mathrm{Adapter}$ [45] - - 0.68 $\times$ 4 99.6 99.6 98.2 96.4 98.5
@@
$\mathrm{OpenVLA}$ TA [21] Vision Backbones 7 $\times$ 4 56.6 58.0 55.6 6.6 44.2
$\mathrm{OpenVLA}$ TA [21] All 7 0.0 0.0 0.0 0.0 0.0
$\mathrm{OpenVLA}$ TA [21] + $\mathbf{S}$ All 7 74.2 82.6 68.8 24.0 62.4
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [21] All 0.68 0.0 0.0 0.0 0.0 0.0
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [21] + $\mathbf{S}$ All 0.68 0.0 0.0 0.0 0.0 0.0
$\mathrm{VLA}$-$\mathrm{Adapter}$ TA [21] + $\mathbf{S}$ Except $\mathbf{H}^{L\rightarrow L}$ 0.70 50.2 34.6 0.0 7.4 23.1
$\mathrm{MergeVLA}_{\mathrm{EMR}}$ EMR [20] 96.0 63.2 62.0 40.6 65.5
$\mathrm{MergeVLA_{TSV}}$ TSV [17] + $\mathbf{S}$ 99.4 97.8 74.4 54.8 81.6
$\mathrm{MergeVLA_{KnOTS}}$ KnOTS[39] + $\mathbf{S}$ 96.8 98.8 84.8 71.4 88.0
$\mathrm{MergeVLA_{TA}}$ TA [21] + $\mathbf{S}$ Except $\mathbf{H}^{L\rightarrow L}$ 0.70 98.0 98.8 85.4 76.6 89.7
$\mathrm{MergeVLA_{WUDI}}$ WUDI [8] + $\mathbf{S}$ 97.6 98.2 85.6 78.2 89.9
$\mathrm{MergeVLA_{TIES}}$ TIES [48] + $\mathbf{S}$ 94.8 94.6 91.8 79.4 90.2
Implementation Details. Our vision-language backbone is Qwen2.5-0.5B [50]. By default, we set $l=L$, $k_{r}=8$, mask ratio $\lambda=0.6$, merging scaling factor $\alpha=1$. All finetuning are conducted on a single NVIDIA A6000 Ada GPU (48 GB). Other hyperparameters can be found in Appendix.
@@ 5.2 Results on LIBERO
We first reproduce results of OpenVLA [28] and VLA-Adapter [45] on four tasks, as highlighted in grey in Table 1. OpenVLA performs notably worse than the others, with an average success rate of 76.5% and only 53.7% on the long-horizon LIBERO-Long task. Our MergeVLA, modified from VLA-Adapter to be merge-friendly, achieves comparable fine-tuning performance, indicating that our structural changes preserve the original capability.
We have tried evaluating single-task finetuned models on unseen tasks (e.g., testing a Spatial expert on the Object suite), all methods achieve 0% success across tasks. This clearly indicates that existing VLA models lack cross-skill generalization. The lower part of Table 1 reports the merging results. Directly applying TA [21] to OpenVLA fails entirely, while merging only the vision and projector components improves to 44.2%. Adding a task-specific mask further raises the average success rate to 62.4%, though performance remains unbalanced. On VLA-Adapter, tight coupling between the VLM and action expert makes merging difficult—even with masking, unless the final action block is excluded. As shown in the blue-highlighted rows, our MergeVLA consistently outperforms other merging baselines. With TIES [48] and WUDI [8], it achieves balanced results across all four tasks (up to 90.2% average success rate), while remaining lightweight and only 6.5% below fine-tuning performance. These results confirm that MergeVLA enables efficient knowledge reuse and strong multi-task performance with a compact VLA model.
@@
$\mathrm{OpenVLA}$ [28] 34.8 0.8 23.0 8.1 28.5 3.5 15.2 16.3
$\mathrm{\pi_{0}}$ [3] 81.4 13.8 58.8 85.0 68.9 6.9 79.0 56.3
$\mathrm{VLA}$-$\mathrm{Adapter}$ [45] 76.6 36.4 73.8 71.0 70.2 37.4 57.2 59.0
$\mathrm{MergeVLA}$ 92.7 62.4 75.7 92.7 73.7 46.4 74.7 72.4
@@
$\mathrm{VLA}$-$\mathrm{Adapter}$ [45] 15.7 6.6 17.6 11.2 15.0 4.1 7.1 10.8
$\mathrm{MergeVLA}_{\mathrm{TSV}}$ 64.3 44.8 58.8 72.9 59.4 30.4 52.7 53.5
$\mathrm{MergeVLA}_{\mathrm{TA}}$ 78.2 53.1 68.4 79.0 65.0 34.6 62.7 61.6
$\mathrm{MergeVLA}_{\mathrm{TIES}}$ 85.7 50.7 66.0 84.2 68.1 30.3 66.0 62.5
@@ 5.3 Results on LIBERO-Plus
From Table 2, we first observe that when using single-task finetuned checkpoints evaluated on their own tasks, MergeVLA exhibits stronger robustness under various perturbations than existing VLA models. This improvement stems from its architecture design, which preserves the pretrained VLM’s inherent robustness to visual and language perturbations. Secondly, in the model-merging setting, we find that applying different merging methods with MergeVLA maintains similar robustness even under cross-task evaluation. Notably, when using TA and TIES merging, the merged model even surpasses OpenVLA, $\mathrm{\pi_{0}}$, and VLA-Adapter in the single-task setting, demonstrating that our MergeVLA can effectively transfer and preserve robustness across tasks.
@@ 6 Conclusion
7 Acknowledgments
This work was partially supported by ARC DE240100105, DP240101814, DP230101196, BA24006, and ARC Industrial Transformation Research Hubs IH230100013.
@@ 9 Algorithm Details
1:Inputs: Task masks $\{\mathbf{S}_{m}\}_{m=1}^{M}$; Expert heads $\{\mathbf{H}_{m}^{\,l\rightarrow L}\}_{m=1}^{M}$; Pretrained VLM weights $\Theta_{0}$; Merged task vector $\tau_{\mathrm{merge}}$; Value projections of the action expert at block $l$: $\mathbf{V}_{\mathrm{T},m}^{\,l},\;\mathbf{V}_{\mathrm{A},m}^{\,l}$; Initial observation $(\mathbf{I}_{0}^{v},\mathbf{I}_{0}^{w},L)$
@@
5:   $\mathbf{h}_{\mathrm{T},m}^{\,l},\;\mathbf{h}_{\mathrm{A},m}^{\,l}=\Theta_{\mathrm{VLM}}^{(m)}(\mathbf{I}_{0}^{v},\mathbf{I}_{0}^{w},L)$ $\triangleright$ Extract $l$-th block hidden states
6:   $\mathbf{V}_{\mathrm{T},m}^{\,l}=\mathbf{L}_{\mathrm{T},m}^{\,l}\,\mathbf{\Sigma}_{\mathrm{T},m}^{\,l}\,(\mathbf{R}_{\mathrm{T},m}^{\,l})^{\top}$
7:   $\mathbf{V}_{\mathrm{A},m}^{\,l}=\mathbf{L}_{\mathrm{A},m}^{\,l}\,\mathbf{\Sigma}_{\mathrm{A},m}^{\,l}\,(\mathbf{R}_{\mathrm{A},m}^{\,l})^{\top}$
8:   $\mathbf{r}_{\mathrm{T},m}=\big\|\mathbf{P}_{\mathrm{T},m}^{\,l}\mathbf{h}^{\,l}_{\mathrm{A},m}\big\|_{2}$ $\triangleright$ Choose top-$r_{k}$ singular vectors from $\mathbf{R}$ to get $\mathbf{P}$
9:   $\mathbf{r}_{\mathrm{A},m}=\big\|\mathbf{P}_{\mathrm{A},m}^{\,l}\mathbf{h}^{\,l}_{\mathrm{T},m}\big\|_{2}$
@@ 10 Preliminary Investigation on OpenVLA
In the early stage of this work, we explored the feasibility of directly applying existing model merging methods to OpenVLA [28], a popular VLA model. OpenVLA consists of three main components: a vision backbone, a projector, and a language model. The language model itself contains 32 transformer blocks followed by a single-layer MLP head (lm_head). We first attempted to merge all components of OpenVLA using Weighted Average and Task Arithmetic methods. However, the merged checkpoint completely failed on all tasks. This was surprising, as OpenVLA is essentially a VLM, and previous studies [13, 6, 37] have shown that VLMs can usually be merged successfully. This prompted us to investigate which part of OpenVLA prevents successful merging.
@@ 10.1 Non-mergeable Components in OpenVLA
To locate the source of failure, we decomposed the model into four submodules: $\mathrm{A.}$ the vision backbone; $\mathrm{B.}$ the projector; $\mathrm{C.}$ the language model body (excluding lm_head); and $\mathrm{D.}$ the lm_head. We then merged each submodule separately across the four official LIBERO task checkpoints using the existing merging method Iso-CTS [31]. Each merged checkpoint was evaluated on 50 trials per subtask. The results are summarized in Table 7, where the gray-highlighted row denotes the single-task fine-tuning performance. From the table, we observe that merging modules $\mathrm{A}$, $\mathrm{B}$, or $\mathrm{D}$ only slightly decreases success rates, whereas merging the language model body ($\mathrm{C}$) causes complete failure on all tasks. This clearly indicates that the language model is the primary source of merging failure.
@@
Table 7: Success rates on the four LIBERO task suites when merging different components of OpenVLA [28] using the Iso-CTS [31] merging method. Each merged checkpoint combines four task-specific models, while unmerged components retain their original weights. During evaluation, each subtask is tested with 50 trials. Gray-highlighted row indicates the success rates of individually fine-tuned models.
@@ 10.2 Progressive Block-wise Merging of the Language Model
Figure 8: Success rate on the LIBERO-Spatial task when progressively merging the first $k$ language model blocks of OpenVLA [28] using the Iso-CTS [31] merging algorithm. Each configuration merges four task-specific checkpoints and is evaluated over 10 trials per subtask.
@@
Figure 9: Mask active ratios for each LIBERO task suite, computed for both the vision backbone and the language model components following the same definition as in the main paper. Masks are obtained using the Task Arithmetic [21] merging method with $\lambda=0.6$.